OpenStudy (anonymous):

The graph of the equation 2x=y^2 from A(0,0) to B(2,2) is rotated around the x axis. The surface area of the resulting solid is?

OpenStudy (anonymous):

Here's what I have so far: http://www.wolframalpha.com/input/?i=y^2%3D2x This needs to be put in terms of x so that, when we integrate we can use dx instead of dy, correct? Also, I had decided that surface area could be found by adding up all the circumferences of the discs from 0 to 2 times dx.

OpenStudy (anonymous):

\[y^2=2x\] turns into \[y=\sqrt(2x)\]

OpenStudy (anonymous):

Assuming the idea of circumferences is correct, C= 2πr and \[\int\limits_{0}^{2}2\pi(\sqrt{2x})dx\]

OpenStudy (turingtest):

your method is not quite right you need ds first

OpenStudy (anonymous):

What is ds?

OpenStudy (turingtest):

I would so it the other way:\[x=\frac{y^2}2\] and\[ds=\sqrt{1+(\frac{dx}{dy})^2}\]

OpenStudy (anonymous):

Uhhh. I don't understand that. maybe I missed something in the lesson...

OpenStudy (turingtest):

correction:\[ds=\sqrt{1+(\frac{dx}{dy})^2}dy\]and the formula for the surface area is then\[S=\int2\pi yds\]

OpenStudy (anonymous):

Yep. I definitely missed something important. thanks for the help. maybe if I review my materials it will look remotely familiar.

OpenStudy (turingtest):

ds is the arc length differential arc length is\[L=\int ds=\int\sqrt{1+(\frac{dy}{dx})}dx\]you need to understand this first before you can understand surface area

OpenStudy (anonymous):

(2*PI)/3 * ((5*5^0.5)-1)

OpenStudy (turingtest):

correction\[L=\int ds=\int\sqrt{1+(\frac{dy}{dx})^2}dx\]

OpenStudy (amistre64):

the latex seems to be crowding up ...

OpenStudy (turingtest):

here's so you can get a feel for arc length http://tutorial.math.lamar.edu/Classes/CalcII/ArcLength.aspx and here is the idea extended to surface area http://tutorial.math.lamar.edu/Classes/CalcII/SurfaceArea.aspx

OpenStudy (turingtest):

@amistre I switched to the SVG math render and that has fixed that problem, as well as a few others

OpenStudy (anonymous):

Thank you.